A 2.4 g plastic bead charged to -3.0 nC and a 4.2
Answer #1

a)

mp = mass of plastic bead = 2.4 x 10-3 kg

qp = charge of plastic bead = - 3 x 10-9 C

mg = mass of glass bead = 4.2 x 10-3 kg

qg = charge of glass bead = 17.1 x 10-9 C

r = distance between beads = 0.02 m

F = force between the beads = k qp qg/r2 = (9 x 109) (3 x 10-9) (17.1 x 10-9)/(0.02)2 = 0.00115 N

ap = acceleration of plastic bead = F/mp = 0.00115/0.0024 = 0.48 m/s2

b)

ag = acceleration of glass bead = F/mg = 0.00115/0.0042 = 0.27 m/s2

c)

towards each other . since they have opposite charge hence force is attaractive between them,